# XiYan-SQL: A Novel Multi-Generator Framework For Text-to-SQL

### v1

@@ IV-A Experimental Setup
Dataset.    We conduct experiments to evaluate the performance of the XiYan-SQL framework on two of the most recognized and challenging benchmarks in Text-to-SQL research in recent years.
BIRD [6], focusing on complex real-world databases, is the most challenging lager-scale cross-domain Text-to-SQL benchmark. It contains 12,751 unique question-SQL pairs and 95 big databases with a total size of 33.4 GB. The training, development, and unpublished test set contain 9,428, 1,534, and 1,789 pairs, respectively. We use its development set for local evaluation and submit our performance results on the unreleased test set for a fair evaluation.
@@
Implementations.    In the proposed framework, all database schemas are represented in the form of the M-schema [35]. For the Schema Filter, we utilize GPT-4o gpt-4o-0806 for all GPT-4o (https://openai.com/index/gpt-4/) [36] as the $\mathcal{M}_{s}$ model for keyword extraction and column selection. We set the maximum number of iterations $p_{s}$ to 2 to obtain the filtered schema set $\mathcal{S}=\{S_{1},S_{2}\}$. For the Multiple SQL Generation part, we first leverage GPT-4o and Qwen-Max https://qwenlm.github.io/zh/blog/qwen2.5-max/ [37] for assistance in constructing relevant multi-task and multi-format SQL data. We conduct experiments on multiple models of varying sizes to thoroughly demonstrate the superiority of our fine-tuning approach. To achieve top performance with XiYan-SQL, we select the widely used Qwen2.5-Coder-32B [38] as the foundation for constructing the final multiple fine-tuned SQL generators. We utilize a total of $p_{m}=5$ generators, which includes one ICL-based generator, ultimately resulting in the generation of $p_{l}=10$ SQL candidates. Due to limited cost and resources, we have not increased the number of models or candidates. For the SQL selection part, we also utilize GPT-4o to augment the training data. Given that the selection task is relatively simpler than the generation task, we choose Qwen2.5-Coder-7B for developing the SQL selection model. All base models used for fine-tuning in the following experiments are their respective instruction versions. For more implementation details, please refer to Appendix II.
@@ IV-B Main Results
BIRD Results. We present the end-to-end performance of the XiYan-SQL on the BIRD benchmark, based on the official benchmark page. In addition to the results obtained from GPT-4o and XiYan-SQL, all other results are derived from the respective original publications. As shown in Table I, our method achieves an EX score of 75.63%, outperforming all previous works and leading the second-best result by 0.84%, establishing a new SOTA performance. This provides strong evidence for the effectiveness of XiYan-SQL as an advanced technique for multiple candidates generation and selection.
@@
Spider Results. To further evaluate the generalizability of XiYan-SQL, we conduct experiments on the Spider test set, with results shown in Table II. Similar to the results obtained on the BIRD benchmark, XiYan-SQL achieves an EX score of 89.65%, surpassing previous methods and establishing a new SOTA performance. It is noteworthy that the testing on Spider does not involve any specific adjustments; instead, we directly employ the model and configuration strategies evaluated on BIRD. This highlights the robust generalizability of our approach, demonstrating consistent performance across various Text-to-SQL scenarios.
@@ IV-D Study of Schema Filter
We first evaluate the performance of our Schema Filter module in filtering database schemas. Results of different schema filter methods on BIRD are shown in the Table IV. We employ the evaluation metrics presented in schema linking [22], where $P$ represents the column’s precision in the filtered schema, while $R_{c}$ and $R_{v}$ denote the recall for columns and values, respectively. Compared to the advanced methods for schema linking that we reproduce, our Schema Filter achieves higher recall for both columns and values while maintaining precision. Furthermore, as iterations increase, it effectively adjusts the precision and recall of the schema, resulting in diverse schemas.
TABLE IV: Schema Filter comparison of different Text-to-SQL methods on BIRD dev.
Methods $P$(%) $R_{c}$(%) $R_{v}$(%)
@@
We further supplement the performance of the final generated SQL queries under different schema linking methods. Due to performance differences among the various generators, we average the results from three different models: the fine-tuned model based on Qwen2.5-Coder-32B, GPT-4o, and Gemini-1.5-pro §§§https://deepmind.google/technologies/gemini/. Notably, we emphasize the $\Delta$EX metrics, which indicate the improvements achieved compared to the full schema method. As shown in Table V, our methods demonstrate significant enhancements. Additionally, we find that selecting a single iteration (i.e., $p_{s}=1$) yields better results than the second iteration. Throughout the entire pipeline, we introduce greater schema diversity through multiple iterations. For additional experiments about the Schema Filter, please refer to Appendix III.
@@ IV-F Study of Multiple SQL Generation
Figure 4 (b) illustrates the performance of our multi-generator ensemble approach varying with the number of candidates. The results for different numbers of candidates are obtained through multiple random samplings from the candidate pool of five SQL generators. As the number of candidates increases, the candidate space exhibits greater diversity, leading to a higher upper-bound; at $p_{l}=10$, this upper-bound reaches 82.2%, showing significant potential. However, the lower-bound also decreases considerably, which indicates that the selection process becomes crucial. Therefore, we explore an enhanced SQL selection method tailored for scenarios involving multiple candidates.
@@ IV-G Study of SQL Selection
After obtaining multiple candidate SQL queries, we introduce a strategy that combines candidate reorganization with a SQL selection model to choose the optimal candidate SQL. Next, we explore the selection strategy with a detailed comparative analysis. Table VIII presents the analysis of different SQL selection methods. We observe that the majority voting strategy, based on consistency, achieves the EX of 69.82%. Furthermore, leveraging our fine-tuned selection model, we compare several distinct selection strategies. Here, $\mathrm{s}_{1}$ denotes presenting the candidate samples to the selection model in a random order, $\mathrm{s}_{2}$ indicates the organization of candidate SQL samples in descending order based on their corresponding model performance, and $\mathrm{s}_{3}$ employs the proposed candidate reorganization to enhance the model’s attention. Our selection strategy (i.e., $\mathrm{s}_{3}$) outperforms the majority voting method by approximately 3% and demonstrated significantly better performance than the other selection strategies. As the number of candidates increases, the selection task becomes more complex, and the advantages of our current method become even more pronounced.
@@
In the aforementioned experiments, we have demonstrated that our selection model, based on the proposed candidate reorganization strategy, achieves performance that significantly surpasses that of the Majority Voting method. This improvement can be attributed not only to our selection strategy, but also to the strong discriminative capability of our selection model. We further evaluate the performance of the selection model. To minimize the impact of candidate quantity and order on the model, we randomly sample three different candidate quantities (2, 6, and 10) from the BIRD dev dataset and organize the candidates in random order for presentation to the model. As shown in Table IX, our fine-tuned selection model outperforms the baseline model and also exceeds the performance of advanced out-of-the-box models (such as GPT-4o and Gemini). Additionally, due to the smaller size of our fine-tuned model, it exhibits faster response times and improved instruction-following capabilities.
@@ IV-H Analysis of Results from Different SQL Generators
In this section, we present the results of different generators that contribute to the final performance of XiYan-SQL. As previously mentioned, we ultimately utilize five distinct generators, with each generator generating two candidates corresponding to two schemas. We provide the individual evaluation results of all the different generators in Table X, along with their contributions to overall performance. Specifically, the generator $\mathrm{SQLG_{1}}$ is a fine-tuned model based on multi-task training without any special formatting; the generator $\mathrm{SQLG_{2}}$ is a fine-tuned model that incorporates multi-task data with complex writing patterns formatting; the generator $\mathrm{SQLG_{3}}$ is a fine-tuned model utilizing multi-task data with standardized presentation styles. The variations in writing patterns and presentation styles are illustrated in Figure 3. The generator $\mathrm{SQLG_{4}}$ is a fine-tuned model that integrates all multi-task data along with various formatted data; the generator $\mathrm{SQLG_{5}}$ is based on the multiple examples ICL method using the GPT-4o model.
For clarity, “Avg. EX” here represents the average of the two generations produced by a single generator based on the two schemas. The “CR” (contribution ratio) refers to the proportion of candidates from each generator after removing all consistent candidates (which account for approximately 45%, where the shortest candidate query is selected). It is essential to note that this ratio ultimately derives from the output of the subsequent SQL selection process.
As observed, the generator $\mathrm{SQLG_{1}}$ exhibits the best performance and contributes the most significantly, representing the optimal performance of a single fine-tuned generation model within our multi-task strategy. The generator $\mathrm{SQLG_{4}}$, serving as a comprehensive mixed data model, also exhibits commendable performance. The generators $\mathrm{SQLG_{2}}$ and $\mathrm{SQLG_{3}}$ focus on diversifying generation pathways by introducing new format data. Meanwhile, the generator $\mathrm{SQLG_{5}}$ leverages the ICL capabilities of GPT-4o to further enhance the upper bound of the candidate pool.
We further present the distribution of the results of each SQL generator within the overall candidate pool, as shown in Table XI. Here, “$R_{n}$” refers to the ratio of the number of correct queries generated by a specific generator to the total number of correct samples when only $n$ generators generate correct queries across all samples. For example, the $R_{1}$ value of $\mathrm{SQLG_{1}}$ is 1.42%, meaning that 1.42% of all correct results are uniquely generated by this generator. As the model with the highest accuracy, $\mathrm{SQLG_{1}}$ exhibits relatively high values of $R_{3}$ and $R_{4}$, suggesting that its generation is more robust and overall quality is higher. In contrast, $\mathrm{SQLG_{3}}$ and $\mathrm{SQLG_{4}}$ show a more concentrated distribution in the middle range, allowing them to explore diverse generations while maintaining quality. On the other hand, $\mathrm{SQLG_{2}}$ and $\mathrm{SQLG_{5}}$ demonstrate higher $R_{1}$ values, indicating greater diversity, and are also capable of generating some unique SQL queries that yield the correct answers.
Furthermore, these two tables also illustrate that, compared to the generator $\mathrm{SQLG_{1}}$ with the highest accuracy, the other fine-tuned generators sacrifice a certain level of precision while introducing diversity. This represents the trade-off between quality and diversity inherent in multiple candidate methods. In summary, each generator plays an important role, and their performances along with their contributions to the candidate pool align with our design objectives for the multi-generator framework of XiYan-SQL. Due to limited resources, further investigation into additional training tasks and methods remains to be explored in future research.
@@
TABLE X: Results of different individual generators in achieving the final performance of XiYan-SQL on the BIRD dev. These results are directly derived from the final pipeline and maintain all configurations.
@@
TABLE XI: The distribution of correct results from different SQL generators at various levels on Bird dev. The $R_{n}$ values, ranging from $R_{1}$ to $R_{5}$, represent the situation where only $n$ generators have generated correct outputs.
@@ IV-I Analysis of Cost Efficiency
During the inference stage, Table XII summarizes the inference costs of XiYan-SQL. It is important to note that the above results reflect the average per-sample statistics of XiYan-SQL on the BIRD dev set. The SQL generation process may include certain self-refine operations, which have been accounted for. We utilize the official API to call GPT-4o, while all other fine-tuned models are deployed using minimal resources. Additionally, due to factors such as network conditions and deployment environments, the results may exhibit some normal fluctuations. In the Schema Filter module, most of the time cost during the retrieval phase arises from the model calls for keyword extraction and the multi-path similarity retrieval, while the time costs during the selection phase result from two iterations of column selection. During the multiple SQL generation phase, the main time consumption involves generating multiple candidate SQL queries, whereas the SQL selection phase only requires the selection of the final result. Ultimately, the end-to-end inference latency for generating 10 candidate SQL queries is approximately 40.5 seconds.
TABLE XII: The average inference costs of the XiYan-SQL framework on BIRD dev.
@@
Schema Filter (Retrieval) GPT-4o 413.9/37 11.8
Schema Filter (Selection) GPT-4o 3791.5/63.1 6.8
Multiple SQL Generation Qwen2.5-Coder-32B 582.5/47.8 2.3
Multiple SQL Generation GPT-4o 763.2/54.7 2.8
SQL Selection Qwen2.5-Coder-7B 1052.3/1 0.8
All pipeline Multiple models 6700/205 40.5

### v2

@@ IV-A Experimental Setup
Dataset.  We evaluate the XiYan-SQL framework on two of the most recognized and challenging benchmarks in modern Text-to-SQL research: BIRD and Spider.
BIRD [13], focusing on complex real-world databases, is the most challenging large-scale cross-domain Text-to-SQL benchmark. It contains 12,751 unique question-SQL pairs and 95 big databases with a total size of 33.4 GB. The training, development, and unpublished test set contain 9,428, 1,534, and 1,789 pairs, respectively. We use its development set for local evaluation and submit our performance results on the unreleased test set for a fair evaluation. Additionally, we utilize the provided Oracle knowledge as evidence.
@@
Implementations.  In the proposed framework, all database schemas are represented in the form of the M-schema [6]. For the Schema Filter, we utilize GPT-4o 333gpt-4o-0806 for all GPT-4o (https://openai.com/index/gpt-4/) [21] as the $\mathcal{M}_{s}$ model for keyword extraction and column selection. We set the maximum number of iterations $p_{s}$ to 2 to obtain the filtered schema set $\mathcal{S}=\{S_{1},S_{2}\}$. For the Multiple SQL Generation part, we first leverage GPT-4o and Qwen-Max  [^4] https://qwenlm.github.io/zh/blog/qwen2.5-max/ [30] for assistance in constructing relevant multi-task and multi-format SQL data. We conduct experiments on multiple models of varying sizes to thoroughly demonstrate the superiority of our fine-tuning approach. To achieve top performance with XiYan-SQL, we select the widely used Qwen2.5-Coder-32B [7] as the foundation for constructing the final multiple fine-tuned SQL generators. We utilize a total of $p_{m}=5$ generators. As detailed in Section IV-H, these consist of four specialized fine-tuned models—a base multi-task model, models for structural and stylistic variations, and a mixed-data model—and one ICL-based generator ($\mathrm{SQLG_{5}}$), ultimately resulting in the generation of $p_{l}=10$ SQL candidates. Due to limited cost and resources, we have not increased the number of models or candidates. For the SQL selection part, we also utilize GPT-4o to augment the training data. Given that the selection task is relatively simpler than the generation task, we choose Qwen2.5-Coder-7B for developing the SQL selection model. All base models used for fine-tuning in the following experiments are their respective instruction versions. For more implementation details, please refer to Appendix II.
@@ IV-B Main Results
BIRD Results. We present the end-to-end performance of the XiYan-SQL on the BIRD benchmark, based on the official benchmark page. In addition to the results obtained from GPT-4o and XiYan-SQL, all other results are derived from the respective original publications. As shown in Table I, our method achieves an EX score of 75.63%. This performance established a new SOTA on the highly competitive BIRD leaderboard at the time of submission (Dec 17, 2024), surpassing all previous works and outperforming the second-place entry by 0.84%. This provides strong evidence for the effectiveness of our “society of experts” philosophy, validating our integrated approach to candidate generation and selection.
@@
Spider Results. To further evaluate the generalizability of XiYan-SQL, we conduct experiments on the Spider test set, with results shown in Table II. Similar to the results obtained on the BIRD benchmark, XiYan-SQL achieves an EX score of 89.65%, surpassing previous methods and establishing a new SOTA performance. This result, achieved without any Spider-specific tuning, demonstrates the framework’s strong performance consistency. It shows that our approach is effective not only on the value-heavy complexity of BIRD but also on the structural complexity emphasized by Spider, complementing our primary findings.
@@ IV-D Study of Schema Filter
We first evaluate the performance of our Schema Filter module in filtering database schemas. Our module iteratively generates a set of diverse schemas. In our experiments, we use the first two schemas generated, denoted as $S_{1}$ (from the first iteration) and $S_{2}$ (from the second iteration). Results for these, alongside different schema filter methods on BIRD, are shown in Table IV. We employ the evaluation metrics presented in schema linking [19], where $P$ represents the column’s precision in the filtered schema, while $R_{c}$ and $R_{v}$ denote the recall for columns and values, respectively. Compared to the advanced methods for schema linking that we reproduce, our Schema Filter achieves higher recall for both columns and values while maintaining precision. Furthermore, as iterations increase, it effectively adjusts the precision and recall of the schema, resulting in diverse schemas.
TABLE IV: Schema Filter comparison of different Text-to-SQL methods on BIRD dev. Our Schema Filter (schema $S_{1}$) and (schema $S_{2}$) refer to the schemas generated in the first and second iterations of our filter, respectively, showcasing the trade-off between precision and recall.
Methods PP(%) $R_{c}$(%) $R_{v}$(%)
@@
[…]
@@ IV-F Study of Multiple SQL Generation
Figure 4 (b) illustrates the performance of our multi-generator ensemble approach as a function of the number of candidates. As these candidates are sampled from our pool of distinct generators, this analysis serves as a proxy for evaluating the impact of the number of generators. The results for different numbers of candidates are obtained through multiple random samplings from the candidate pool of five SQL generators. As the number of candidates increases, the candidate space exhibits greater diversity, leading to a higher upper-bound; at $p_{l}=10$, this upper-bound reaches 82.2%, showing significant potential. However, the lower-bound also decreases considerably, which indicates that the selection process becomes crucial. Therefore, we explore an enhanced SQL selection method tailored for scenarios involving multiple candidates. For a more direct and comprehensive analysis of the interaction between the number of generators and the Schema Filter iterations, please refer to the detailed study in Appendix IV.A.
In summary, the experiments in this section demonstrate that a multi-generator approach significantly outperforms single-generator, as well as simple sampling or multi-prompt methods. The success of this paradigm relies on both the quantity and, crucially, the diversity of the generators. While the multi-task strategy (Section IV-E) enhances the quality of each individual generator, it is our multi-format training strategy that is specifically designed to induce the structural and stylistic diversity across the ensemble. To rigorously quantify the individual and synergistic contributions of these two strategies, a comprehensive, decoupled ablation study is provided in Appendix IV.B. The results therein confirm that the quality-enhancing multi-task strategy and the diversity-inducing multi-format strategy provide complementary benefits, validating our overall fine-tuning methodology.
@@ IV-G Study of SQL Selection
After obtaining multiple candidate SQL queries, we introduce a strategy that combines candidate reorganization with a SQL selection model to choose the optimal candidate SQL. Next, we explore the selection strategy with a detailed comparative analysis. Table VIII presents the analysis of different SQL selection methods. We observe that the majority voting strategy, based on consistency, achieves the EX of 69.82%. Furthermore, leveraging our fine-tuned selection model, we compare several distinct selection strategies. Here, $\mathrm{s}_{1}$ denotes presenting the candidate samples to the selection model in a random order, $\mathrm{s}_{2}$ indicates the organization of candidate SQL samples in descending order based on their corresponding model performance, and $\mathrm{s}_{3}$ employs the proposed candidate reorganization to enhance the model’s attention. Our selection strategy (i.e., $\mathrm{s}_{3}$) outperforms the majority voting method by approximately 3% and demonstrates significantly better performance than the other selection strategies. As the number of candidates increases, the selection task becomes more complex, and the advantages of our current method become even more pronounced.
@@
In the aforementioned experiments, we have demonstrated that our selection model, based on the proposed candidate reorganization strategy, achieves performance that significantly surpasses that of the Majority Voting method. This improvement highlights the powerful synergy between our selection strategy and the strong discriminative capability of our model. The strategy effectively amplifies the model’s ability to focus on the correct SQL. We further evaluate the performance of the selection model. To minimize the impact of candidate quantity and order on the model, we randomly sample three different candidate quantities (2, 6, and 10) from the BIRD dev dataset and organize the candidates in random order for presentation to the model. As shown in Table IX, our fine-tuned selection model outperforms the baseline model and also exceeds the performance of advanced out-of-the-box models (such as GPT-4o and Gemini). Additionally, due to the smaller size of our fine-tuned model, it exhibits faster response times and improved instruction-following capabilities.
@@ IV-H Analysis of Results from Different SQL Generators
In this section, we present the results of different generators that contribute to the final performance of XiYan-SQL. As previously mentioned, we ultimately utilize five distinct generators, with each generator generating two candidates corresponding to two schemas. We provide the individual evaluation results in Table X, along with their contributions to overall performance. Specifically, the roles of generators are as follows: the generator $\mathrm{SQLG_{1}}$ is a base fine-tuned model based on multi-task training; $\mathrm{SQLG_{2}}$ is a specialized model fine-tuned on a dataset enriched with structural variations (e.g., CTEs); $\mathrm{SQLG_{3}}$ is a specialized model fine-tuned on a dataset featuring specific stylistic variations; $\mathrm{SQLG_{4}}$ is a comprehensive model fine-tuned on a mix of all multi-task and multi-format data; $\mathrm{SQLG_{5}}$ is an ICL-based generator using the GPT-4o model.
Avg. EX represents the average execution accuracy of the two candidates from a single generator. The CR (Contribution Ratio) is defined to measure the contribution of each generator in more difficult cases that require careful selection. The calculation is as follows: we first identify and exclude the “unanimous” cases (approx. 45% of the dataset) where all SQL candidates yield identical execution results. For the remaining cases, our SQL selection method chooses a single optimal candidate. The CR for a generator is the percentage of these non-unanimous cases where the selected candidate is generated by that generator. Since each selected candidate belongs to a single generator, the sum of all CRs is 100%.
As observed, the generator $\mathrm{SQLG_{1}}$ exhibits the best individual performance (Avg. EX) and contributes the most significantly (CR), representing the strong capability of our multi-task fine-tuning strategy. Notably, $\mathrm{SQLG_{2}}$, which is trained to favor complex “structural variations”, exhibits a lower Avg. EX score than $\mathrm{SQLG_{1}}$. This suggests that generating these advanced forms is inherently harder and more error-prone. A similar principle of trade-offs applies to $\mathrm{SQLG_{3}}$ and $\mathrm{SQLG_{4}}$; their respective focuses on strict stylistic consistency and broad solution-space coverage come at the cost of a minor drop in peak accuracy on standard patterns. However, its non-trivial Contribution Ratio and its unique correct generations (shown by the $R_{1}$ value in Table XI) demonstrate that it is essential for solving certain challenging queries that simpler forms cannot handle. These findings validate our core hypothesis: no single generation form is universally optimal; instead, integrating diverse generation pathways is a highly effective strategy for tackling a wide spectrum of real-world challenges.
The analysis of the result distribution in Table XI further supports this, where the $R_{n}$ values denote the situation in which only $n$ generators have generated correct outputs, as explained in the table note. As the model with the highest accuracy, $\mathrm{SQLG_{1}}$ exhibits high values of $R_{3}$ and $R_{4}$, suggesting that its generation is more robust and overall quality is higher. In contrast, the distributions of $\mathrm{SQLG_{3}}$ and $\mathrm{SQLG_{4}}$ are more concentrated within the middle range, exploring diverse generations while maintaining quality. Meanwhile, $\mathrm{SQLG_{2}}$ and $\mathrm{SQLG_{5}}$ exhibit higher $R_{1}$ values, indicating greater diversity; each possesses a unique ability to solve certain problems alone. This illustrates the trade-off between quality and diversity inherent in our ensemble approach.
Case-Level Analysis We present a detailed case-level analysis from the BIRD dev set as qualitative evidence for the value of structural variation. The full details are shown in Figure 5. This question requires a two-step reasoning process: first, identifying the county with the most closed schools, and second, listing all schools within that specific county. Our best-performing single model, $\mathrm{SQLG_{1}}$, failed to solve this. It generated a query with a flawed aggregation strategy, attempting to solve both reasoning steps at once. In contrast, the specialized model $\mathrm{SQLG_{2}}$ successfully resolved the query by generating a correct and more readable solution using a CTE, which effectively decomposed the problem into two distinct logical steps. This case powerfully illustrates our core argument. The structural diversity introduced by $\mathrm{SQLG_{2}}$—specifically, its ability to generate a CTE—is not merely for variety. It is a critical component for achieving correctness on challenging queries that require sophisticated logical decomposition. While $\mathrm{SQLG_{2}}$ is not entirely generalizable, it achieves important complementarity with other models.
Figure 5: Case study demonstrating the value of Structural Variation on a challenging query from the BIRD dev. Inputs such as evidence and the schema are omitted for clarity.
In summary, each generator plays an important and distinct role, and their performances align with our design objectives for the multi-generator framework of XiYan-SQL.
@@
TABLE X: Performance of the five individual generators on the BIRD dev set. These results are directly derived from the final pipeline and maintain all configurations.
@@
TABLE XI: Distribution of correct results across different levels of inter-generator agreement on the BIRD dev. The column $R_{n}$ shows the percentage of correct answers from a generator in cases where exactly $n$ models succeeded. Thus, $R_{1}$ measures the contribution of unique solutions, while $R_{5}$ represents universally correct answers.
@@ IV-I Analysis of Cost Efficiency
During the inference stage, Table XII summarizes the inference costs of XiYan-SQL. It is important to note that the above results reflect the average per-sample statistics of XiYan-SQL on the BIRD dev set. The SQL generation process may include certain self-refine operations, which have been accounted for. We utilize the official API to call GPT-4o, while all other fine-tuned models are deployed using minimal resources. Additionally, due to factors such as network conditions and deployment environments, the results may exhibit some normal fluctuations. In the Schema Filter module, most of the time cost during the retrieval phase arises from the model calls for keyword extraction and the multi-path similarity retrieval, while the time costs during the selection phase result from two iterations of column selection. During the multiple SQL generation phase, the main time consumption involves generating multiple candidate SQL queries. The SQL Selection stage includes two parts: executing all candidates against the database and inferring the selection model to select the final result. On our test environment (Apple M4 Pro, 48GB RAM), the average execution time per sample (with 10 candidates) on BIRD dev is 0.35 seconds. It is important to note that this execution time is highly dependent on the specific hardware and the scale of the database. Ultimately, the total end-to-end latency per sample is approximately 40.81 seconds.
To further illustrate the inference cost, we compare XiYan-SQL with the previous SOTA method, CHASE-SQL[22]. The performance achieved by our method with 10 generated SQL candidates surpasses that of CHASE-SQL, which generates 21 SQL candidates. In this configuration, based on the inference information provided by CHASE-SQL, the overall latency of our pipeline is only 60% of that of CHASE-SQL.
TABLE XII: The average inference costs of the XiYan-SQL framework on BIRD dev. *SQL execution latency is measured on a local machine (Apple M4 Pro, 48 GB RAM) and can vary depending on the hardware and database size.
@@
Schema Filter (Retrieval) GPT-4o 413.9/37 11.75
Schema Filter (Selection) GPT-4o 3791.5/63.1 6.77
Multiple SQL Generation (per candidate, avg.) Qwen2.5-Coder-32B 582.5/47.8 2.25
Multiple SQL Generation (per candidate, avg.) GPT-4o 763.2/54.7 2.75
Candidate SQL Execution N/A (Database) N/A $0.35^{*}$
SQL Selection Qwen2.5-Coder-7B 1052.3/1 0.78
All pipeline Multiple models $\sim\!6700/\sim\!205$ $\sim\!40.81$